# Ameliorating transient noise bursts in gravitational-wave searches for intermediate-mass black holes

### v1

@@ Abstract
The direct observation of intermediate-mass black holes (IMBH) populations would not only strengthen the possible evolutionary link between stellar and supermassive black holes, but unveil the details of the pair-instability mechanism and elucidate their influence in galaxy formation. Conclusive observation of IMBHs remained elusive until the detection of gravitational-wave (GW) signal GW190521, which lies with high confidence in the mass gap predicted by the pair-instability mechanism. Despite falling in the sensitivity band of current GW detectors, IMBH searches are challenging due to their similarity to transient bursts of detector noise, known as glitches. In this proof-of-concept work, we combine a matched-filter algorithm with a Machine Learning (ML) method to differentiate IMBH signals from non-transient burst noise, known as glitches. In particular, we build a multi-layer perceptron network to perform a multi-class classification of the output triggers of matched-filter. In this way we are able to distinguish simulated GW IMBH signals from different classes of glitches that occurred during the third observing run (O3) in single detector data. We train, validate, and test our model on O3a data, reaching a true positive rate of over $90\%$ for simulated IMBH signals. In O3b, the true positive rate is over $70\%$. We also combine data from multiple detectors to search for simulated IMBH signals in real detector noise, providing a significance measure for the output of our ML method.
@@ I.1 Motivation
As stellar evolution models predict, stars with helium core masses in the range $\sim 32-64\,\text{M}_{\odot}$ are subject to pulsation pair instability, while stars with helium core masses in the range $\gtrsim 50-130\,\text{M}_{\odot}$ leave no remnant due to pair-instability supernovae (PISN) Woosley et al. (2007); Woosley and Heger (2021). Nonetheless, we find intermediate-mass black holes (IMBH) within this mass gap and beyond, spawning from $\sim 10^{2}-10^{5}\,\text{M}_{\odot}$.
@@
Although Advanced LIGO Aasi et al. (2015a) and Advanced Virgo Acernese et al. (2015) gravitational-wave (GW) detectors detected 11 candidates during the first observing run (O1) and the second observing run (O2) Abbott et al. (2019a, b, 2021a), the detection of IMBHs in GW searches remained elusive until the detection of GW190521 during the third observing run (O3) Abbott et al. (2021b, 2022). The estimate of the individual source-frame mass components of GW190521 was $(m_{1},m_{2})=(85^{+11}_{-14},66^{+17}_{18})\,\text{M}_{\odot}$, with a final remnant mass of $M_{f}=142^{+28}_{-16}\,\text{M}_{\odot}$ making it the first conclusive direct observation of an IMBH Abbott et al. (2020). Final analyses of O3 data yielded even higher mass IMBH GW events. In particular, GW190426_190642 Abbott et al. (2021a) supersedes GW190521 with a final mass of $M_{f}=172.9^{+37.7}_{-33.6}\,\text{M}_{\odot}$, while GW200220_061928 Abbott et al. (2023) has probably the highest final mass of the second half of O3 (O3b) $M_{f}=141^{+51}_{-31}\,\text{M}_{\odot}$.
@@
The origin of supermassive black holes (SMBHs) of masses beyond $10^{5}\,\text{M}_{\odot}$ remains a fundamental mystery, despite their presence in nearly every galaxy, including the Milky Way Richstone et al. (1998); Collaboration (2022). Recent observational surveys with the James Webb Telescope Gardner et al. (2006) challenge models of SMBH formation and growth, raising questions about the origin of seed black holes and the mechanisms enabling their rapid growth to supermassive scales. The PISN mass gap implies that stellar collapse alone may not account for SMBH formation. A plausible scenario is the hierarchical mergers of IMBHs, providing a potential evolutionary bridge between stellar-mass black holes and SMBHs. Greene et al. (2020); Abbott et al. (2022). Directly observing IMBH populations would not only strengthen the proposed evolutionary link and unveil the role of the PISN mechanism, but also serve as a rigorous test of General Relativity, given the characteristically strong merger and ringdown signals they produce Isi et al. (2019).
@@
Despite falling in the sensitivity band of current GW interferometers, GW signals from IMBH mergers are challenging as few cycles of the signal can be observed with current ground-based detectors. Initial, GW IMBH searches were restricted to using the model waveform-independent algorithm coherent Wave Burst (cWB)Klimenko et al. (2005, 2006, 2011) and a ringdown templated search Aasi et al. (2015b) in order to probe the merger-ringdown phase. The improvement of detector sensitivity at low frequencies in the advanced era allowed to probe the short inspiral phase with matched-filtering techniques. Nowadays, state-of-the-art searches employ the modeled-independent cWB in its IMBH configuration Klimenko et al. (2016); Gayathri et al. (2019); Drago et al. (2021), and matched-filter-based techniques using modeled-waveforms Allen et al. (2012); Messick et al. (2017); Usman et al. (2016); Sachdev et al. (2019); Chandra et al. (2021a, b); Tsukada et al. (2023).
@@
Despite the improvement in sensitivity, IMBH searches are still hampered by non-Gaussian transient burst noise, known as glitches, that can mask or mimic GW candidates, reduce the amount of analyzable data increasing the noise floor and affect the estimation of the power spectral density (PSD) Abbott et al. (2018a); Acernese et al. (2023); Glanzer et al. (2023). As an example, in Fig. 1 we show the similarities between a Blip glitch and an IMBH signal. Blips are short glitches (0.20.20.2\,0.2s) of unknown origin that have a characteristic morphology of a symmetric ‘teardrop’ shape in time-frequency in the range [30,250] Hz. As we can observe in Fig. 1, Blips have a similar duration to IMBH signals and intersect the frequency band of these signals.
@@
Figure 1: (Left) Q-transform of a Blip glitch retrieved from Gravity Spy (Zevin et al., 2017). (Right) Q-transform of a gravitational wave (GW) signal from the merger of two high mass black holes to form an IMBH with total mass $106.6^{+13.5}_{-14.8}M_{\odot}$.
@@ I.2 Previous work
Through the years, scientists have implemented multivariate ML methods to enhance conventional searches, allowing them to more accurately distinguish GW signals from background glitches. The implemented ML methods would learn a single feature vector derived from multi-detector analysis to perform the binary classification task. Such applications have proven successful in the context of binary black holes Baker et al. (2015); Hodge (2014), gamma-ray bursts Adams et al. (2013); Kim et al. (2015), and burst searches Mishra et al. (2021); Szczepańczyk et al. (2023); Gayathri et al. (2020); Lopez et al. (2022a). Nonetheless, these multivariate methods are not restricted to a single feature vector; they can also integrate information from several single-detector analysis Kapadia et al. (2017) and other data representations, such as singular value decomposition  Magee et al. (2024) to distinguish GW signals from background glitches.
@@ I.3 Relation to previous work
It may be possible to construct a more comprehensive data structure that fully captures the unique features separating GW signals from glitches. While earlier approaches rely on a single feature vector, our method instead utilizes sets of feature vectors derived from matched-filtering. This process correlates the input detector data with collections of modeled GW signals—referred to as templates—which are organized into template banks.
The key idea is that when we “match” the unknown detector data $s(t)$ with the templates of a bank, a handful of templates, say $N$, will be sufficiently similar to $s(t)$ within a finite time window $\Delta t$. Associated to each matching template, the search pipeline will launch several “triggers”, all flagging the same potential candidate. We refer to this inherent cluster structure as cluster track, or simply tracks, inspired by the tracking produced in particle colliders Shiltsev and Zimmermann (2020). Given the morphological time-frequency differences between glitches and GW signals, we expect these “clusters of triggers” to exhibit a meaningful structure to differentiate them with ML methods. We do so by introducing a new statistic: a multivariate multi-class probability vector, which measures the probability that each track is generated by either an astrophysical signal or by one of the glitch classes considered.
In this work we use the tracks generated in single-detector, right after the matched-filtering step. Hence, to maintain as much information as possible, we reproduce the IMBH search of GstLAL during O3 using LIGO Hanford (H1), LIGO Livingston (L1) and Virgo (V1) Abbott et al. (2022). With this information, the ML algorithm learns to distinguish glitches from simulated IMBH signals in a controlled environment. Afterwards, we use the algorithm to detect a set of simulated IMBH, performing a background estimation. This proof-of-concept study demonstrates that even with a restricted template bank, it is possible to effectively separate signals from glitches, making a potential impact on the accuracy of GW analysis. It is relevant to to note that while this particular investigation focuses on IMBH, this method can be extended to other compact binary coalescence (CBC) signals.
@@ II Matched Filtering pipeline
As the typical amplitude of a GW signal $h(t)$ is orders of magnitude smaller than the detector background noise $n(t)$, sophisticated detection algorithms, or pipelines, are needed to identify CBC signals. Most of them are based on matched-filter, which in turn relies on the precise models of the sources, for instance, employing Post-Newtonian approaches Blanchet et al. (1995) or the effective-one-body formalism Buonanno and Damour (1999). Some examples are the GstLAL-based inspiral Messick et al. (2017); Sachdev et al. (2019); Hanna et al. (2020a); Cannon et al. (2021); Tsukada et al. (2023), PyCBC Allen et al. (2012); Allen (2005); Canton et al. (2014); Usman et al. (2016); Nitz et al. (2017); Davies et al. (2020); Nitz et al. (2018), MBTA Adams et al. (2016); Aubin et al. (2021) and SPIIR Chu et al. (2022) pipelines. In this work, we use the GstLAL-based inspiral pipeline, hereafter refered to as GstLAL, so in the following sections we provide a high-level overview of some of its particularities.
@@ II.1 Time-domain matched filtering
The tasks of a matched-filter search is to find a target signal $h(t)$ given the detector output $s(t)=n(t)+h(t)$. This is achieved by cross-correlating $s(t)$ with a modelled waveform $h_{m}(t)$, known as a template. We can define a time-dependent complex scalar product as Allen et al. (2012),
[…]
where $S_{n}(f)$ represents the one-sided PSD, and $\tilde{\cdot}$ and denote the Fourier transform and the complex conjugate, respectively. We can also define the real and imaginary part of the scalar product as ${\langle s|h\rangle(t)=(s|h)(t)+i(s|h)(t)}$, where we define $(s|h)(t)=\Re\langle s|h\rangle(t)$. With this definition and given $h(t)$, its normalized time series is $\hat{h}=h/(h|h)$.
The GstLAL analysis performs matched filtering directly in the time domain Sachdev et al. (2019); Messick et al. (2017). In accordance with the notation of Messick et al. (2017), and given a time series $s(t)$ and a normalized complex template $h_{c}(t)$, the output of the matched-filter technique is a complex time series,
@@ II.2 Template banks
CBC template waveforms depend on the intrinsic parameters of the source $\lambda_{int}$, such as the masses ($m_{1},m_{2}$) and z-spins ($s_{1z},s_{2z}$) of the binary components, and on extrinsic parameters $\lambda_{ext}$, which are related to the position of the source concerning the observer, such as the luminosity distance $D$. Thus, the goal of matched filtering is to maximize the detection statistic, SNR, over all these parameters. Luckily, for aligned-spin signals, the effect of the extrinsic parameters can be absorbed into a constant scale factor and a phase Creighton and Anderson (2011) and the SNR (Eq. 3) maximizes over such extrinsic parameters. Therefore, we are only left with the maximisation over the intrinsic parameters which is usually performed by a brute force approach, where the SNR time-series is computed for a dense grid of modelled GW signals, called templates, characterized by $\lambda_{int}$. Templates are gathered in large template banks, which are usually generated Ajith et al. (2014) to achieve a minimal loss of SNR due to the discreteness of the grid, while keeping a manageable computational cost.
Different pipelines employ different strategies to cross-correlate template banks with the output data from the detector while mitigating glitches. In this work we are interested in utilizing the product of this cross-correlation targeting IMBH signals, which are heavily harmed by glitches due to their short duration and similar frequency range. For this aim, we employ GstLAL to reproduce the IMBH search of O3 Abbott et al. (2022). We use the IMBH template bank which uses $44902$ templates to cover aligned-spin systems with total mass $M=m_{1}+m_{2}$ in range $[50,600]\,\text{M}_{\odot}$ and mass ratio $q=\frac{m_{1}}{m_{2}}$ between $[1,10]$, with both spins in the range $[-0.98,0.98]$. The starting frequency of the analysis is at 101010\,10Hz. For further details the interested reader can refer to Abbott et al. (2022).
@@ II.3 $\xi^{2}$-statistic
The detector strain contains glitches that can mask or mimic GW signals, producing large peaks in the SNR time series. As glitches in the detector could yield large SNRs, to mitigate them GstLAL also calculates a signal consistency check, known as $\xi^{2}$-statistic, whenever it records an SNR above a certain threshold. This check is performed by determining the similarity between the SNR time series of the data and the expected SNR time series from the real signal within a $\delta t$ time window around the peak Messick et al. (2017). Mathematically, the $\xi^{2}$-statistic is defined as,
[…]

### v2

@@ Abstract
The direct observation of intermediate-mass black holes (IMBH) populations would not only strengthen the possible evolutionary link between stellar and supermassive black holes, but unveil the details of the pair-instability mechanism and elucidate their influence in galaxy formation. Conclusive observation of IMBHs remained elusive until the detection of gravitational-wave (GW) signal GW190521, which lies with high confidence in the mass gap predicted by the pair-instability mechanism. Despite falling in the sensitivity band of current GW detectors, IMBH searches are challenging due to their similarity to transient bursts of detector noise, known as glitches. In this proof-of-concept work, we combine a matched-filter algorithm with a Machine Learning (ML) method to differentiate IMBH signals from non-transient burst noise, known as glitches. In particular, we build a multi-layer perceptron network to perform a multi-class classification of the output triggers of matched-filter. In this way we are able to distinguish simulated GW IMBH signals from different classes of glitches that occurred during the third observing run (O3) in single detector data. We train, validate, and test our model on O3a data, reaching a true positive rate of over $90\%$ for simulated IMBH signals. To test the generalization ability over the evolutionary observing run, we test on the useen data of O3b, which yields a true positive rate of over $70\%$ . We also combine data from multiple detectors to search for simulated IMBH signals in real detector noise, providing a significance measure for the output of our ML method.
@@ I.1 Motivation
As stellar evolution models predict, stars with helium core masses in the range $\sim 32-64\,\text{M}_{\odot}$ are subject to pulsation pair instability, while stars with helium core masses in the range $\gtrsim 50-130\,\text{M}_{\odot}$ leave no remnant due to pair-instability supernovae (PISN) Woosley:2007qp; Woosley:2021xba. Nonetheless, we find intermediate-mass black holes (IMBH) within this mass gap and beyond, spawning from $\sim 10^{2}-10^{5}\,\text{M}_{\odot}$.
@@
Although Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos gravitational-wave (GW) detectors detected 11 candidates during the first observing run (O1) and the second observing run (O2) LIGOScientific:2018mvr; LIGOScientific:2019ysc; LIGOScientific:2021usb, the detection of IMBHs in GW searches remained elusive until the detection of GW190521 during the third observing run (O3) LIGOScientific:2020ibl; LIGOScientific:2021tfm. The estimate of the individual source-frame mass components of GW190521 was $(m_{1},m_{2})=(85^{+11}_{-14},66^{+17}_{18})\,\text{M}_{\odot}$, with a final remnant mass of $M_{f}=142^{+28}_{-16}\,\text{M}_{\odot}$ making it the first conclusive direct observation of an IMBH PhysRevLett.125.101102. Final analyses of O3 data yielded even higher mass IMBH GW events. In particular, GW190426_190642 LIGOScientific:2021usb supersedes GW190521 with a final mass of $M_{f}=172.9^{+37.7}_{-33.6}\,\text{M}_{\odot}$, while GW200220_061928 LIGOScientific:2021djp has probably the highest final mass of the second half of O3 (O3b) $M_{f}=141^{+51}_{-31}\,\text{M}_{\odot}$.
@@
The origin of supermassive black holes (SMBHs) of masses beyond $10^{5}\,\text{M}_{\odot}$ remains a fundamental mystery, despite their presence in nearly every galaxy, including the Milky Way Richstone:1998ky; EHCSgtA*. Recent observational surveys with the James Webb Telescope Gardner:2006ky challenge models of SMBH formation and growth, raising questions about the origin of seed black holes and the mechanisms enabling their rapid growth to supermassive scales. The PISN mass gap implies that stellar collapse alone may not account for SMBH formation. A plausible scenario is the hierarchical mergers of IMBHs, providing a potential evolutionary bridge between stellar-mass black holes and SMBHs. 2020ARA&A..58..257G; LIGOScientific:2021tfm. Directly observing IMBH populations would not only strengthen the proposed evolutionary link and unveil the role of the PISN mechanism, but also serve as a rigorous test of General Relativity, given the characteristically strong merger and ringdown signals they produce Isi2019HierarchicalTO.
@@
Despite falling in the sensitivity band of current GW interferometers, GW signals from IMBH mergers are challenging as few cycles of the signal can be observed with current ground-based detectors. Initial, GW IMBH searches were restricted to using the model waveform-independent algorithm coherent Wave Burst (cWB)PhysRevD.72.122002; SKlimenko_2006; PhysRevD.83.102001 and a ringdown templated search Aasi_2015 in order to probe the merger-ringdown phase. The improvement of detector sensitivity at low frequencies in the advanced era allowed to probe the short inspiral phase with matched-filtering techniques. Nowadays, state-of-the-art searches employ the modeled-independent cWB in its IMBH configuration Klimenko:2015ypf; Gayathri:2019omo; Drago:2020kic, and matched-filter-based techniques using modeled-waveforms Allen:2005fk; Messick:2016aqy; usman2016pycbc; Sachdev:2019vvd; Chandra:2021wbw; Chandra:2021xvs; Tsukada:2023edh.
@@
Despite the improvement in sensitivity, IMBH searches are still hampered by non-Gaussian transient burst noise, known as glitches, that can mask or mimic GW candidates, reduce the amount of analyzable data increasing the noise floor and affect the estimation of the power spectral density (PSD) LIGOScientific:2017tza; Virgo:2022ysc; Glanzer:2022avx. As an example, in Fig. 1 we show the similarities between a Blip glitch and an IMBH signal. Blips are short glitches (0.20.2\,s) of unknown origin that have a characteristic morphology of a symmetric ‘teardrop’ shape in time-frequency in the range [30,250] Hz. As we can observe in Fig. 1, Blips have a similar duration to IMBH signals and intersect the frequency band of these signals.
@@
Figure 1: (Left) Q-transform of a Blip glitch retrieved from Gravity Spy (zevin2017gravity). (Right) Q-transform of a gravitational wave (GW) signal from the merger of two high mass black holes to form an IMBH with total mass $106.6^{+13.5}_{-14.8}M_{\odot}$.
@@ I.2 Previous work
Through the years, scientists have implemented multivariate ML methods to enhance conventional searches, allowing them to more accurately distinguish GW signals from background glitches. The implemented ML methods would learn a single feature vector derived from multi-detector analysis to perform the binary classification task. Such applications have proven successful in the context of binary black holes Baker:2014eba; Hodge:2014yca, gamma-ray bursts Adams:2013pna; Kim:2014nba, and burst searches Mishra:2021tmu; Szczepanczyk:2022urr; Gayathri:2020bly; Lopez:2021ikt. Nonetheless, these multivariate methods are not restricted to a single feature vector; they can also integrate information from several single-detector analysis Kapadia:2017fhb and other data representations, such as singular value decomposition  2024arXiv241015513M to distinguish GW signals from background glitches.
@@ I.3 Relation to previous work
It may be possible to construct a more comprehensive data structure that fully captures the unique features separating GW signals from glitches. While earlier approaches rely on a single feature vector Baker:2014eba; Hodge:2014yca; Adams:2013pna; Kim:2014nba; Mishra:2021tmu; Szczepanczyk:2022urr; Gayathri:2020bly; Lopez:2021ikt, our method instead utilizes sets of feature vectors derived from matched-filtering. This process correlates the input detector data with collections of modeled GW signals—referred to as templates—which are organized into template banks.
The key idea is that when we “match” the unknown detector data $s(t)$ with the templates of a bank, a handful of templates, say $N$, will be sufficiently similar to $s(t)$ within a finite time window $\Delta t$. Associated to each matching template, the search pipeline will launch several “triggers”, all flagging the same potential candidate. We refer to this inherent cluster structure as cluster track, or simply tracks, inspired by the tracking produced in particle colliders Shiltsev2020ModernAF. Given the morphological time-frequency differences between glitches and GW signals, we expect these “clusters of triggers” to exhibit a meaningful structure to differentiate them with ML methods. We do so by introducing a new statistic: a multivariate multi-class probability vector, which measures the probability that each track is generated by either an astrophysical signal or by one of the glitch classes considered.
In this work we use the tracks generated in single-detector, right after the matched-filtering step. Hence, to maintain as much information as possible, we reproduce the IMBH search of GstLAL during O3 using LIGO Hanford (H1), LIGO Livingston (L1) and Virgo (V1) LIGOScientific:2021tfm. With this information, the ML algorithm learns to distinguish glitches from simulated IMBH signals in a controlled environment. Afterwards, we use the algorithm to detect a set of simulated IMBH, performing a background estimation. This proof-of-concept study demonstrates that even with a restricted template bank, it is possible to effectively separate signals from glitches, making a potential impact on the accuracy of GW analysis. It is relevant to to note that while this particular investigation focuses on IMBH, this method can be extended to other compact binary coalescence (CBC) signals.
@@ II Matched Filtering pipeline
As the typical amplitude of a GW signal $h(t)$ is orders of magnitude smaller than the detector background noise $n(t)$, sophisticated detection algorithms, or pipelines, are needed to identify CBC signals. Most of them are based on matched-filter, which in turn relies on the precise models of the sources, for instance, employing Post-Newtonian approaches PhysRevLett.74.3515 or the effective-one-body formalism Buonanno:1998gg. Some examples are the GstLAL-based inspiral 2017Messick; Sachdev:2019vvd; hanna2020fast; cannon2021gstlal; Tsukada:2023edh, PyCBC Allen:2005fk; allen2005chi; dal2014implementing; usman2016pycbc; nitz2017detecting; davies2020extending; nitz2018rapid, MBTA adams2016low; aubin2021mbta and SPIIR chu2022spiir pipelines. In this work, we use the GstLAL-based inspiral pipeline, hereafter refered to as GstLAL, so in the following sections we provide a high-level overview of some of its particularities.
@@ II.1 Time-domain matched filtering
The tasks of a matched-filter search is to find a target signal $h(t)$ given the detector output $s(t)=n(t)+h(t)$. This is achieved by cross-correlating $s(t)$ with a modelled waveform $h_{m}(t)$, known as a template. We can define a time-dependent complex scalar product as Allen:2005fk,
$\langle s|h\rangle(t)=2\int_{-\infty}^{\infty}\frac{\tilde{h}_{m}^{*}(f)\tilde{s}(f)}{S_{n}(f)}e^{i2\pi ft}df,$ (1)
where $S_{n}(f)$ represents the one-sided PSD, and $\tilde{\cdot}$ and denote the Fourier transform and the complex conjugate, respectively. We can also define the real and imaginary part of the scalar product as $\langle s|h\rangle(t)=(s|h)(t)+i[s|h](t)$}, where we define $(s|h)(t)=\Re\langle s|h\rangle(t)$} and $s|h](t)=\Im\langle s|h\rangle(t)$}. With this definition and given $h(t)$, its normalized time series is $\hat{h}=h/(h|h)$.
The GstLAL analysis performs matched filtering directly in the time domain Sachdev:2019vvd; 2017Messick. In accordance with the notation of 2017Messick, and given a time series $s(t)$ and a normalized complex template $h_{c}(t)$, the output of the matched-filter technique is a complex time series,
@@ II.2 Template banks
CBC template waveforms depend on the intrinsic parameters of the source $\lambda_{int}$, such as the masses ($m_{1},m_{2}$) and z-spins ($s_{1z},s_{2z}$) of the binary components, and on extrinsic parameters $\lambda_{ext}$, which are related to the position of the source concerning the observer, such as the luminosity distance $D$. Thus, the goal of matched filtering is to maximize the detection statistic, SNR, over all these parameters. Luckily, for aligned-spin signals, the effect of the extrinsic parameters can be absorbed into a constant scale factor and a phase Creighton:2011zz and the SNR (Eq. 3) maximizes over such extrinsic parameters. Therefore, we are only left with the maximisation over the intrinsic parameters which is usually performed by a brute force approach, where the SNR time-series is computed for a dense grid of modelled GW signals, called templates, characterized by $\lambda_{int}$. Templates are gathered in large template banks, which are usually generated Ajith:2012mn to balance between a small loss of SNR due to the discreteness of the grid and a manageable computational cost.
Different pipelines employ different strategies to cross-correlate template banks with the output data from the detector while mitigating glitches. In this work we are interested in utilizing the product of this cross-correlation targeting IMBH signals, which are heavily harmed by glitches due to their short duration and similar frequency range. For this aim, we employ GstLAL to reproduce the IMBH search of O3 LIGOScientific:2021tfm. We use the IMBH template bank which uses $44902$ templates to cover aligned-spin systems with total mass $M=m_{1}+m_{2}$ in range $50,600]\,\text{M}_{\odot}$ and mass ratio $q=\frac{m_{1}}{m_{2}}$ between $[1,10]$, with both spins in the range [0.98,0.98][-0.98,0.98]. The starting frequency of the analysis is at 1010\,Hz. For further details the interested reader can refer to LIGOScientific:2021tfm.
@@ II.3 $\xi^{2}$-statistic
The detector strain contains glitches that can mask or mimic GW signals, producing large peaks in the SNR time series. As glitches in the detector could yield large SNRs, to mitigate them GstLAL also calculates a signal consistency check, known as $\xi^{2}$-statistic, whenever it records an SNR above a certain threshold. This check is performed by determining the similarity between the SNR time series of the data and the expected SNR time series from the real signal within a $\delta t$ time window around the peak 2017Messick. Mathematically, the $\xi^{2}$-statistic is defined as,
$\xi^{2}=\frac{\int_{-\delta t}^{\delta t}|z(t)-z(0)R(t)|^{2}dt}{\int_{-\delta t}^{\delta t}(2-2|R(t)|^{2})dt},$ (4)